OpenStudy (anonymous):

Solve x divided by negative six is greater than three for x. x is greater than negative two x is less than negative eighteen x is greater than negative eighteen x is less than negative two PLZ PLZ HELP ASAP

OpenStudy (anonymous):

I replied to your previous question about this :)

OpenStudy (anonymous):

i probably didnt see it let me check

OpenStudy (anonymous):

yes! its that i need the answer

OpenStudy (anonymous):

Well lets do it here. \[\frac{x}{-6}>3=\frac{-x}{6}>3\] To simplify, isolate x first. Remember, dividing both sides by a negative number flips the inequality sign.

OpenStudy (anonymous):

Ok the = sign in between is not practical :) I meant the inequality on the left is equal to the inequality on the right

OpenStudy (anonymous):

so x is greater than 3?

OpenStudy (anonymous):

Lets do it in two steps. First, multiply both sides by 6: \[\frac{-x}{6}>3\] \[-x>18\]

OpenStudy (anonymous):

Now, x is not nicely isolated yet, as it has a minus sign in front of it. Multiplying both sides by -1 is the logical next step. Remember though, multiplying both sides by a negative number flips the > sign.

OpenStudy (anonymous):

Could you continue from \[-x > 18?\]

OpenStudy (anonymous):

Or is it clear? :)